Q2. Use Cramers Rule to find the solution of the following system: 21 +02 = 3 -3.01 +203 = 0 22 - 2.13 = 2.
Answer #1

-2 312 GIVEN Soli- System is x + xq =3 - 3%, + 2X, = 0 769 - De z Matrix farm is its o 3 76 -3 &2 1 Compare it with AX we hav3 now, Dx, 0 2 1 a ) Dai 3 (0-2) -1 (0-4) 2 - 6+4 Da, Do 2 Dag 5 2 -2 -) D? 1(0-4) - 3 (6-0) -4-18 & Da=-22 Ox, -3 T 2 Dag al- x ) i; ㅋ x = 2. -8. - l ㅑ 2분 X mlos ) which ů the stequined delo